OpenStudy (anonymous):

2. M(8, 7) is the midpoint of . The coordinates of S are (9, 5). What are the coordinates of R? (16, 14) (8.5, 6) (10, 3) (7, 9) 3.Find the distance between points P(8, 2) and Q(3, 8) to the nearest tenth. 14.9 61 7.8 11 4.The Frostburg-Truth bus travels on a straight road from Frostburg Mall to Sojourner Truth Park. The mall is 4 miles east and 4 miles north of the City Center. The park is 2 miles west and 4 miles south of the Center. How far is it from the mall to the park to the nearest tenth of a mile? 11.1 miles 10.0 miles 5.7 miles 4.5 miles

OpenStudy (anonymous):

@johncooper

OpenStudy (anonymous):

first question: take your midpoint's x and y and multiply by two then subtract S's x and y. you'll then have R's coordinates

OpenStudy (anonymous):

IM SO CONFUSED

OpenStudy (anonymous):

second question: the distance formula is \[d \sqrt{\left( x_{2}-x _{1} \right)^{2}+\left( y _{2}-y _{1} \right)^{2}} \]

OpenStudy (anonymous):

for the first question what are you confused about?

OpenStudy (anonymous):

i dont get the part of if im supposed to multiply 8and7 together or seperate

OpenStudy (anonymous):

seperate

OpenStudy (anonymous):

for ex

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

8x2=x x-9=?

OpenStudy (anonymous):

that would be the x coordinate of R

OpenStudy (anonymous):

ohh ok thanks

OpenStudy (anonymous):

and the third problem: The city center is (0,0) on a graph. the mall is (4,4) and the park is (-2,-4). use the distance formula just like the second problem to find the answer

OpenStudy (anonymous):

for the 2nd one i got 25+12=37 which i know is not right

OpenStudy (anonymous):

alright lets work on that one

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

so you know the distance formula

OpenStudy (anonymous):

now take (8,2) first

OpenStudy (anonymous):

put it into the formula

OpenStudy (anonymous):

having 8 be x1 and 2 being y1

OpenStudy (anonymous):

what next

OpenStudy (anonymous):

then take the other problem and plug it into the same way except let 3 be x2 and 8 be y2

OpenStudy (anonymous):

\[d \sqrt{\left( 3-8 \right)^{2}\left( 8-2 \right)^{2}}\] this is what you should have

OpenStudy (anonymous):

i got it

OpenStudy (anonymous):

so what is your answer?

OpenStudy (anonymous):

\[5^2+6^2\]

OpenStudy (anonymous):

it should be a -5

OpenStudy (anonymous):

61

OpenStudy (anonymous):

and remember to find the square root of that answer

OpenStudy (anonymous):

i got it wrong the 1st time because i accidently multiplyed 6and 2

OpenStudy (anonymous):

so what is your final answer?

OpenStudy (anonymous):

why find the squ.........7.8

OpenStudy (anonymous):

yep good job

OpenStudy (anonymous):

and just do the same thing on the last question

OpenStudy (anonymous):

i got another ?

OpenStudy (anonymous):

what is it?

OpenStudy (anonymous):

the third problem has 3 equations why if im going to do the distance formula

OpenStudy (anonymous):

just do the distance from the park to the mall. ignore city central.

OpenStudy (anonymous):

thanks for like the 100th time

OpenStudy (anonymous):

nope problem